# The Contiguous Art Gallery Problem is in $\Theta(n\log n)$

### v2

@@ 2. $\mathbf{B}$
Let $e\in E$ be the edge defining $\ell$. By definition, $e$ lies on $\ell$ and thus $\ell$ separates $e$ from $i$ and $j$. Consider the directed line $\ell_{j}$ from the right endpoint of $e$ to $j$, and consider the chain from $e$’s right endpoint to $j$. This chain must contain at least one edge $e^{*}$ whose start lies left of $\ell_{j}$ and whose endpoint right right of $\ell_{j}$ (recall that we distinguish between right and strictly right). It follows that $\mathcal{E}[u,v]$ lies left of $\ell_{j}$. Defining $\ell_{i}$ analogously yields a wedge bounded by $\ell_{i}$ and $\ell_{j}$ that contains $\mathcal{E}[u,v]$ (see the blue region in Figure˜6). The fact that one of $u$ or $v$ is in $H_{i,j}$ implies that the point of intersection between $\ell_{i}$ and $\ell_{j}$ lies in $H_{g}$ and not on $t$. Moreover, $g$ also lies in $H_{g}$ and not on $t$, thus no point on $t$ is in $\mathcal{E}[u,v]$, which contradicts $g^{*}$ being in $\mathcal{E}[u,v]$. ∎
@@ 4.2  Bad and ugly dominators
Definition 5 (Bad dominators ($\star$)).
@@
Finally, the ugly dominators are defined a bit differently as they are defined off of an existing bad guard $(g,[u,v])$:
Definition 7 (Ugly dominators (➷)).
@@
Figure 8: Illustration of Theorem˜7 which shows a bad contiguous guard $(g,[u,v])$.
@@
Proof.
Let now $S_{v}$ be the shortest path from $u$ to $v$ and denote by $\ell_{u}$ the directed supporting line of its first edge. By Lemma˜3, $S_{v}$ forms a left-turning chain. Let $S$ be the shortest path from $u$ to $j+1$. Since $(g,[u,v])$ has maximal $v$, Lemma˜2 states that $\overline{g\,v}$ contains a reflex vertex of $P$. If $v\neq j+1$, $\overline{g\,v}$ contains a reflex vertex in its interior. This implies the first edge of $S_{v}$ and $S$ are the same, and thus the line $\ell_{u}$ equals $\ell$. If $v=j+1$, then $S_{v}$ is equal to $S$, so also $\ell$ equals $\ell_{u}$. Denote by $P^{\prime}$ the convex polygon that is $\mathcal{E}[i-1,j+1]$ intersected with the half-plane right of $\ell_{u}$, see Figure˜8. Denote by $c$ a point in $P^{\prime}$ that sees farthest along $[j,j+1]$. We can choose $c$ to be a vertex of $P^{\prime}$. We make a case distinction:
@@
    if $c$ is not a vertex of $\mathcal{E}[i-1,j+1]$ then $c$ must be an intersection point of $\ell_{u}=\ell$ and $\mathcal{E}[i-1,j+1]$. As $\ell$ and $\mathcal{E}[i-1,j+1]$ intersect, the ugly dominator $(g^{*},[u,v^{*}])$ exists, where $g^{*}$ is the last point of intersection between $\ell$ and $\mathcal{E}[i-1,j+1]$. As the chain $S_{v}$ is a left-turning chain, the intersection point of $\ell$ and $\mathcal{E}[i-1,j+1]$ that sees furthest along the edge $[j,j+1]$ is the last point of intersection, i.e. $g^{*}=c^{\prime}$. As $c^{\prime}$ sees at least as far as $g$, because $g\in P^{\prime}$, it must be that $v^{*}\geq v$. ∎
@@ 4.3  The next function and our high-level algorithm
Definition 8 (The next function [theEnemy, RobsonSpaldingZheng2024_AnalyticArcCover]).
@@
We show how to transform $G$ into such a set $G^{\prime}$ (see also Figure˜8). We start by replacing every guard $(g,[u,v])\in G$ by a guard $(g,[u^{\prime},v^{\prime}])$ such that $u,v]\subset[u^{\prime},v^{\prime}]$ and the interval $u^{\prime},v^{\prime}]$ is maximal to form $G^{\prime}$. Next, we show that we can adapt $G^{\prime}$ such that either there is a good guard in $G^{\prime}$ or a guard whose visibility ends at a vertex of $P$. Lemma˜5 implies that this good guard is dominated by a guard in $\mathcal{D}$, concluding the proof.
@@
Suppose that $G^{\prime}$ contains only bad guards and let $(g,[u,v])\in G^{\prime}$ be such a bad guard. If $(u,v)$ does not contain a vertex of $P$ then $u,v]\subset[i,i+1]$ for some vertex $i$. We now replace $(g,[u,v])$ by $(i,[i,i+1])$ and conclude the proof. Similarly, if $v$ is a vertex of $P$ then, because $|G|>1$, there has to be at least one other guard $(g^{\prime},[u^{\prime},v^{\prime}])\in G^{\prime}$ such that $v\in[u^{\prime},v^{\prime}]$. We replace this guard by $(g^{\prime},[v,v^{\prime}])$ so that the start of its chain is a vertex of $P$ which concludes the proof. Lastly, note that it is not possible for $g$ to be $u$ (resp. $v$), as in that case, $g$ sees all edges of $P$ containing $g=u$ (resp. $g=v$), and in particular, strictly more than $[u,v]$, contradicting the maximality of $[u,v]$.
@@
It follows that for all bad guards $(g,[u,v])\in G^{\prime}$, the open chain $(u,v)$ contains a vertex of $P$, while neither $u$ nor $v$ is a vertex of $P$. By Lemma˜2, the segment $\overline{v\,g}$ contains a reflex vertex $r$ of $P$ with $r\notin[u,v]$. Hence, there exists a guard $(g^{\prime},[u^{\prime},v^{\prime}])$ such that $r\in[u^{\prime},v^{\prime}]$. If $u^{\prime}=r$ or $v^{\prime}=r$, we conclude the proof as above. Otherwise, if $r\in(u^{\prime},v^{\prime})$, then $g^{\prime}$ lies left of both supporting lines of the edges incident to $r$. This implies that $g^{\prime}$ lies left of $\overline{v\,g}$. Since $\sphericalangle(v,g,u)>\pi$, any point that sees $r$ and lies left of $\overline{v\,g}$ must also lie within the polygon $P_{g}$ defined by $[u,v]$, $\overline{v\,g}$, and $\overline{g\,u}$ (see Figure˜8). Observe that the points in $u^{\prime},v^{\prime}]\setminus(u,v)$ form a single closed chain; otherwise, $u,v]\subset[u^{\prime},v^{\prime}]$, contradicting the minimality of $G^{\prime}$. Define $a,b]=[u^{\prime},v^{\prime}]\setminus(u,v)$. If we replace $(g^{\prime},[u^{\prime},v^{\prime}])$ by $(g^{\prime},[a,b])$, the resulting set of guards still covers $\partial P$. Crucially, $[a,b]$ lies left of $\overline{g^{\prime}\,v}$ and left of either the supporting line of $\overline{u\,g^{\prime}}$ or $\overline{g\,g^{\prime}}$. As $g^{\prime}$ lies in $P_{g}$ and $\sphericalangle(v,g,u)>\pi$, this implies that, unless $g^{\prime}=a$ or $g^{\prime}=b$, $\sphericalangle(a,g^{\prime},b)\leq\pi$, and thus $G^{\prime}$ contains a good guard. Finally, if $g^{\prime}=a$ (resp. $g^{\prime}=b$), then $g^{\prime}$ is on $\partial P$ and in particular, sees the entirety of the edge containing $a$ (resp. $b$), hence we can extend $[a,b]$ to start (or end) in a vertex, which concludes the proof. ∎
@@
Proof.
We apply to each $x\in X$ the function $\textnormal{{next}}(x)$ recursively until a set of guards covering $\partial P$ is obtained. Each such sequence contains at most $k+1$ guards and this thus takes $O(k\,T\log n)$ time.
We claim that one of these guard sequences must have size $k$. By Theorem˜7 there exists at least one guarding solution of size $k$ that contains a guard $(g,[x,v])$ where the guard is either in $\mathcal{D}$ or $x$ is a vertex of $P$. A classical argument now implies that recursively applying $\textnormal{{next}}(x)$ finds a solution of size $k$ also. For completeness, we write the argument here: Let $G=\{(g_{i},[v_{i-1},v_{i}])\}$ be minimum ordered sequence of guards that guards $\partial P$ that is generated by recursively applying $\textnormal{{next}}(x)$. Then $(g_{1},[x_{1},x_{2}])=(g,[x,\textnormal{{next}}(x)])$.
Suppose, for the sake of contradiction, that $|G|>k$. Let $G^{*}$ be an optimal solution of size $k$ that includes $(g,[x,v])$ and order this solution along $\partial P$, starting at $(g,[x,v])$. Since $|G|>|G^{*}|$, there must exist a minimum index $i$ such that for the $i$th guards $(g_{i},[v_{i-1},v_{i}])\in G$ and $(g_{i}^{*},[u_{i}^{*},v_{i}^{*}])\in G^{*}$, it holds that $v_{i}<v_{i}^{*}$. However, since $i$ is minimum, it must be that $v_{i-1}\geq u_{i}^{*}$. But then $v_{i-1}\in[u_{i}^{*},v_{i}^{*}]$ which makes $\textnormal{{next}}(v_{i-1})$ return a value that is at least $v_{i}^{*}$—a contradiction. ∎
@@ 5  Intermezzo: data structures and sliding sequences
The second property follows from a standard adaptation of this data structure (see also [Eades2020Visibility]): for each pair of consecutive edges on a precomputed shortest path, determine whether the turn is left or right. For each node in the tree, store a symbol indicating whether all of its descendant edges are left-turning, right-turning, or mixed. For a query, we can determine for each of the $O(\log n)$ newly computed bridges whether the turn at its endpoints is left or right. Combined with the pre-stored symbols, this information suffices to answer the second query. ∎
@@
Throughout the paper, we assume that we have access to ˜1 and ˜2 for the given polygon. We repeatedly use the concept of a sliding sequence. A sliding sequence can be thought of as a pair of indices $(i,j)$ that can be updated by incrementing either $i$ or $j$. Formally we define in Definition˜1 a conforming sliding sequence used by almost all of our algorithms. For each $(i,j)$, we maintain the visibility core $\mathcal{E}[i-1,j+1]$ as a convex-hull data structure supporting emptiness, containment, extreme-point, and ray-shooting queries, and an additional angle query (defined later). For any linear-size sliding sequence $\{(i_{t},j_{t})\}$, Chan, Hershberger, and Pratt [Chan2019Time-Windowed] give a data structure supporting such updates and queries in $O(\log n)$ time:
Lemma 9 (Paraphrased Lemma 1 in [Chan2019Time-Windowed]).
For any linear-size sliding sequence $\sigma$, the total number of distict vertices that get added or removed from the visibility core $\mathcal{E}[i-1,j+1]$ when iterating over $(i,j)\in\sigma$ is linear.
Lemma 10.
@@
Proof.
Let $v$ be a vertex of the visibility core $\mathcal{E}[i-1,j+1]$ that is not a vertex of $\mathcal{E}[i^{\prime}-1,j^{\prime}+1]$, where $(i^{\prime},j^{\prime})$ follows $(i,j)$ in $\sigma$. There are two possible reasons for $v$ not being a vertex of $\mathcal{E}[i^{\prime}-1,j^{\prime}+1]$: either one of its supporting edges has been removed from the sliding window, or a new half-plane has been added to the visibility core that excludes $v$. In the first case, since removed edges never reappear in later windows, $v$ cannot return. In the second case, the new half-plane remains in the sliding window at least as long as the edges defining $v$, again preventing $v$ from reappearing. Hence, each vertex appears in at most one contiguous interval. ∎
@@
  • Increment $i$: Delete the corresponding half-plane.
  • Increment $j$: Insert a new half-plane.
@@
Proof.
Let $(i,j)$ be the dynamic index pair from the Sliding-Window data structure (3). We maintain $\mathcal{E}[i-1,j+1]$ and $\mathcal{E}[i-1,j+2]$ separately, subject to increments of $i$ and $j$ via Corollary˜11. This maintains the vertices of these convex areas as balanced binary trees where the left-to-right traversal of the leaves corresponds to a clockwise traversal. For a a convex polygon with $n$ vertices, whose vertices are stored in a balanced tree according to their cyclic order, the containment, ray shooting, and extreme-point queries can be answered in $O(\log n)$ time [DOBKIN1983]. We can thus immediately support all queries except for angle queries.
For angle queries, conceptually rotate and translate the plane until $\overline{u\,v}$ corresponds with the horizontal line through $(0,0)$ with the area left of $\overline{u\,v}$ lying above this line. For any fixed angle $\gamma>0$, the points $\vec{x}\in\mathbb{R}^{2}_{\geq 0}$ for which $\sphericalangle(u,\vec{x},v)=\gamma$ forms a convex arc. For $\gamma_{1},\gamma_{2}$ with $\gamma_{1}>\gamma_{2}$ the corresponding convex arc lies strictly above the arc induced by $\gamma_{2}$. It follows that the function $\vec{x}\mapsto-\sphericalangle(u,\vec{x},v)$ defined for all points $\vec{x}\in\mathbb{R}^{2}_{\geq 0}$ is well-defined, unimodal in $\mathcal{E}[i-1,j+1]$, and strictly monotone between adjacent edges of $\mathcal{E}[i-1,j+1]$. Since we have the edges of $\mathcal{E}[i-1,j+1]$ explicitly, in-order, in a balanced binary tree, it then follows that the point in $g$ that maximizes $\sphericalangle(i,g,j)$ lies on the boundary of $\mathcal{E}[i-1,j+1]$ and that we can find it via performing binary search along the edges of $\mathcal{E}[i-1,j+1]$ in $O(\log n)$ time. ∎

### v4

@@ 2. $\mathbf{B}$
Let $e\in E$ be the edge defining $\ell$. By definition, $e$ lies on $\ell$ and thus $\ell$ separates $e$ from $i$ and $j$. Consider the directed line $\ell_{j}$ from the right endpoint of $e$ to $j$, and consider the chain from $e$’s right endpoint to $j$. This chain must contain at least one edge $e^{*}$ whose start lies left of $\ell_{j}$ and whose endpoint right right of $\ell_{j}$ (recall that we distinguish between right and strictly right). It follows that $\mathcal{E}[u,v]$ lies left of $\ell_{j}$. Defining $\ell_{i}$ analogously yields a wedge bounded by $\ell_{i}$ and $\ell_{j}$ that contains $\mathcal{E}[u,v]$ (see the blue region in Figure˜6). The fact that one of $u$ or $v$ is in $H_{i,j}$ implies that the point of intersection between $\ell_{i}$ and $\ell_{j}$ lies in $H_{g}$ and not on $t$. Moreover, $g$ also lies in $H_{g}$ and not on $t$, thus no point on $t$ is in $\mathcal{E}[u,v]$, which contradicts $g^{*}$ being in $\mathcal{E}[u,v]$.
@@ 4.1 Bad and ugly dominators
Definition 4.11 (Bad dominators ($\star$)).
@@
Finally, the ugly dominators are defined off of an existing bad guard $(g,[u,v])$:
Definition 4.13 (Ugly dominators (➷)).
@@
Figure 8: Illustration of Theorem˜4.17 which shows a bad contiguous guard $(g,[u,v])$.
@@
Proof 4.15.
Let now $S_{v}$ be the shortest path from $u$ to $v$ and denote by $\ell_{u}$ the directed supporting line of its first edge. By Lemma˜4.4, $S_{v}$ forms a left-turning chain. Let $S$ be the shortest path from $u$ to $j+1$. Since $(g,[u,v])$ has maximal $v$, Lemma˜4.2 states that $\overline{g\,v}$ contains a reflex vertex of $P$. If $v\neq j+1$, $\overline{g\,v}$ contains a reflex vertex in its interior. This implies the first edge of $S_{v}$ and $S$ are the same, and thus the line $\ell_{u}$ equals $\ell$. If $v=j+1$, then $S_{v}$ is equal to $S$, so also $\ell$ equals $\ell_{u}$. Denote by $P^{\prime}$ the convex polygon that is $\mathcal{E}[i-1,j+1]$ intersected with the half-plane right of $\ell_{u}$, see Figure˜8. Denote by $c$ a point in $P^{\prime}$ that sees farthest along $[j,j+1]$. We can choose $c$ to be a vertex of $P^{\prime}$. We make a case distinction:
@@
    if $c$ is not a vertex of $\mathcal{E}[i-1,j+1]$ then $c$ must be an intersection point of $\ell_{u}=\ell$ and $\mathcal{E}[i-1,j+1]$. As $\ell$ and $\mathcal{E}[i-1,j+1]$ intersect, the ugly dominator $(g^{*},[u,v^{*}])$ exists, where $g^{*}$ is the last point of intersection between $\ell$ and $\mathcal{E}[i-1,j+1]$. As the chain $S_{v}$ is a left-turning chain, the intersection point of $\ell$ and $\mathcal{E}[i-1,j+1]$ that sees furthest along the edge $[j,j+1]$ is the last point of intersection, i.e. $g^{*}=c^{\prime}$. As $c^{\prime}$ sees at least as far as $g$, because $g\in P^{\prime}$, it must be that $v^{*}\geq v$.
@@ 4.2 The next function and our high-level algorithm
Definition 4.16 (The next function [theEnemy, RobsonSpaldingZheng2024_AnalyticArcCover]).
@@
We show how to transform $G$ into such a set $G^{\prime}$ (see also Figure˜8). We start by replacing every guard $(g,[u,v])\in G$ by a guard $(g,[u^{\prime},v^{\prime}])$ such that $u,v]\subseteq[u^{\prime},v^{\prime}]$ and the interval $u^{\prime},v^{\prime}]$ is maximal to form $G^{\prime}$. Next, we show that we can adapt $G^{\prime}$ such that either there is a good guard in $G^{\prime}$ or a guard whose visibility ends at a vertex of $P$. Lemma˜4.9 implies that this good guard is dominated by a guard in $\mathcal{D}$, concluding the proof.
@@
Suppose that $G^{\prime}$ contains only bad guards and let $(g,[u,v])\in G^{\prime}$ be such a bad guard. If $(u,v)$ does not contain a vertex of $P$ then $u,v]\subseteq[i,i+1]$ for some vertex $i$. We now replace $(g,[u,v])$ by $(i,[i,i+1])$ and conclude the proof. Similarly, if $v$ is a vertex of $P$ then, because $|G|>1$, there has to be at least one other guard $(g^{\prime},[u^{\prime},v^{\prime}])\in G^{\prime}$ such that $v\in[u^{\prime},v^{\prime}]$. We replace this guard by $(g^{\prime},[v,v^{\prime}])$ so that the start of its chain is a vertex of $P$ which concludes the proof. Lastly, note that it is not possible for $g$ to be $u$ (resp. $v$), as in that case, $g$ sees all edges of $P$ containing $g=u$ (resp. $g=v$), and in particular, strictly more than $[u,v]$, contradicting the maximality of $[u,v]$.
@@
It follows that for all bad guards $(g,[u,v])\in G^{\prime}$, the open chain $(u,v)$ contains a vertex of $P$, while neither $u$ nor $v$ is a vertex of $P$. By Lemma˜4.2, the segment $\overline{v\,g}$ contains a reflex vertex $r$ of $P$ with $r\notin[u,v]$. Hence, there exists a guard $(g^{\prime},[u^{\prime},v^{\prime}])$ such that $r\in[u^{\prime},v^{\prime}]$. If $u^{\prime}=r$ or $v^{\prime}=r$, we conclude the proof as above. Otherwise, if $r\in(u^{\prime},v^{\prime})$, then $g^{\prime}$ lies left of both supporting lines of the edges incident to $r$. This implies that $g^{\prime}$ lies left of $\overline{v\,g}$. Since $\sphericalangle(v,g,u)>\pi$, any point that sees $r$ and lies left of $\overline{v\,g}$ must also lie within the polygon $P_{g}$ defined by $[u,v]$, $\overline{v\,g}$, and $\overline{g\,u}$ (see Figure˜8). Observe that the points in $u^{\prime},v^{\prime}]\setminus(u,v)$ form a single closed chain; otherwise, $u,v]\subseteq[u^{\prime},v^{\prime}]$, contradicting the minimality of $G^{\prime}$. Define $a,b]=[u^{\prime},v^{\prime}]\setminus(u,v)$. If we replace $(g^{\prime},[u^{\prime},v^{\prime}])$ by $(g^{\prime},[a,b])$, the resulting set of guards still covers $\partial P$. Crucially, $[a,b]$ lies left of $\overline{g^{\prime}\,v}$ and left of either the supporting line of $\overline{u\,g^{\prime}}$ or $\overline{g\,g^{\prime}}$. As $g^{\prime}$ lies in $P_{g}$ and $\sphericalangle(v,g,u)>\pi$, this implies that, unless $g^{\prime}=a$ or $g^{\prime}=b$, $\sphericalangle(a,g^{\prime},b)\leq\pi$, and thus $G^{\prime}$ contains a good guard. Finally, if $g^{\prime}=a$ (resp. $g^{\prime}=b$), then $g^{\prime}$ is on $\partial P$ and in particular, sees the entirety of the edge containing $a$ (resp. $b$), hence we can extend $[a,b]$ to start (or end) in a vertex, which concludes the proof.
@@
Proof 4.20.
We apply to each $x\in X$ the function $\textnormal{{next}}(x)$ recursively until a set of guards covering $\partial P$ is obtained. Each such sequence contains at most $k+1$ guards and this thus takes O(kTn)O(k\,Tn) time.
We claim that one of these guard sequences must have size $k$. By Theorem˜4.17 there exists at least one guarding solution of size $k$ that contains a guard $(g,[x,v])$ where the guard is either in $\mathcal{D}$ or $x$ is a vertex of $P$. A classical argument now implies that recursively applying $\textnormal{{next}}(x)$ finds a solution of size $k$ also. For completeness, we write the argument here: Let $G=\{(g_{i},[v_{i-1},v_{i}])\}$ be minimum ordered sequence of guards that guards $\partial P$ that is generated by recursively applying $\textnormal{{next}}(x)$. Then $(g_{1},[x_{1},x_{2}])=(g,[x,\textnormal{{next}}(x)])$.
Suppose, for the sake of contradiction, that $|G|>k$. Let $G^{*}$ be an optimal solution of size $k$ that includes $(g,[x,v])$ and order this solution along $\partial P$, starting at $(g,[x,v])$. Since $|G|>|G^{*}|$, there must exist a minimum index $i$ such that for the $i$th guards $(g_{i},[v_{i-1},v_{i}])\in G$ and $(g_{i}^{*},[u_{i}^{*},v_{i}^{*}])\in G^{*}$, it holds that $v_{i}<v_{i}^{*}$. However, since $i$ is minimum, it must be that $v_{i-1}\geq u_{i}^{*}$. But then $v_{i-1}\in[u_{i}^{*},v_{i}^{*}]$ which makes $\textnormal{{next}}(v_{i-1})$ return a value that is at least $v_{i}^{*}$—a contradiction.
@@ 5 Intermezzo: data structures and sliding sequences
The second property follows from a standard adaptation of this data structure (see also [Eades2020Visibility]): for each pair of consecutive edges on a precomputed shortest path, determine whether the turn is left or right. For each node in the tree, store a symbol indicating whether all of its descendant edges are left-turning, right-turning, or mixed. For a query, we can determine for each of the $O(\log n)$ newly computed bridges whether the turn at its endpoints is left or right. Combined with the pre-stored symbols, this information suffices to answer the second query.
@@
Throughout the paper, we assume that we have access to ˜1 and ˜2 for the given polygon. We repeatedly use the concept of a sliding sequence. A sliding sequence can be thought of as a pair of indices $(i,j)$ that can be updated by incrementing either $i$ or $j$. Formally we define in Definition˜3.1 a conforming sliding sequence used by almost all of our algorithms. For each $(i,j)$, we maintain the visibility core $\mathcal{E}[i-1,j+1]$ as a convex-hull data structure supporting emptiness, containment, extreme-point, and ray-shooting queries, and an additional angle query (defined later). For any linear-size sliding sequence $\{(i_{t},j_{t})\}$, Chan, Hershberger, and Pratt [Chan2019Time-Windowed] give a data structure supporting such updates and queries in $O(\log n)$ time:
Lemma 5.2 (Paraphrased Lemma 1 in [Chan2019Time-Windowed]).
For any linear-size sliding sequence $\sigma$, the total number of distinct vertices that get added or removed from the visibility core $\mathcal{E}[i-1,j+1]$ when iterating over $(i,j)\in\sigma$ is linear.
Lemma 5.3.
@@
Proof 5.4.
Let $v$ be a vertex of the visibility core $\mathcal{E}[i-1,j+1]$ that is not a vertex of $\mathcal{E}[i^{\prime}-1,j^{\prime}+1]$, where $(i^{\prime},j^{\prime})$ follows $(i,j)$ in $\sigma$. There are two possible reasons for $v$ not being a vertex of $\mathcal{E}[i^{\prime}-1,j^{\prime}+1]$: either one of its supporting edges has been removed from the sliding window, or a new half-plane has been added to the visibility core that excludes $v$. In the first case, since removed edges never reappear in later windows, $v$ cannot return. In the second case, the new half-plane remains in the sliding window at least as long as the edges defining $v$, again preventing $v$ from reappearing. Hence, each vertex appears in at most one contiguous interval.
@@
[…]
@@
Proof 5.7.
Let $(i,j)$ be the dynamic index pair from the Sliding-Window data structure (3). We maintain $\mathcal{E}[i-1,j+1]$ and $\mathcal{E}[i-1,j+2]$ separately, subject to increments of $i$ and $j$ via Corollary˜5.5. This maintains the vertices of these convex areas as balanced binary trees where the left-to-right traversal of the leaves corresponds to a clockwise traversal. For a a convex polygon with $n$ vertices, whose vertices are stored in a balanced tree according to their cyclic order, the containment, ray shooting, and extreme-point queries can be answered in $O(\log n)$ time [DOBKIN1983]. We can thus immediately support all queries except for angle queries.
For angle queries, conceptually rotate and translate the plane until $\overline{u\,v}$ corresponds with the horizontal line through $(0,0)$ with the area left of $\overline{u\,v}$ lying above this line. For any fixed angle $\gamma>0$, the points $\vec{x}\in\mathbb{R}^{2}_{\geq 0}$ for which $\sphericalangle(u,\vec{x},v)=\gamma$ forms a convex arc. For $\gamma_{1},\gamma_{2}$ with $\gamma_{1}>\gamma_{2}$ the corresponding convex arc lies strictly above the arc induced by $\gamma_{2}$. It follows that the function $\vec{x}\mapsto-\sphericalangle(u,\vec{x},v)$ defined for all points $\vec{x}\in\mathbb{R}^{2}_{\geq 0}$ is well-defined, unimodal in $\mathcal{E}[i-1,j+1]$, and strictly monotone between adjacent edges of $\mathcal{E}[i-1,j+1]$. Since we have the edges of $\mathcal{E}[i-1,j+1]$ explicitly, in-order, in a balanced binary tree, it then follows that the point in $g$ that maximizes $\sphericalangle(i,g,j)$ lies on the boundary of $\mathcal{E}[i-1,j+1]$ and that we can find it via performing binary search along the edges of $\mathcal{E}[i-1,j+1]$ in $O(\log n)$ time.